Assignment 2 Exercise 21.40 く) 13of15 Constants Part A A-3.50 sC point charge is at the origin, and a second -600 nC point charge is on to Nos at z Find the electric field (magnitude and direction) at point on the z-axis at r -0200 m Express your answer with the appropriate units. Enter positive value if the field is in the positive z-direction and negative value if the field is in the negative a-direction. 525 Submit Previous Amwers Request Answer X Incorrect; Try Again: 3 attempts remaining
Part B Find the electric field (magnitude and direction) at point on the z-axis at r 1.20 m. Express your answer with the appropriate units. Enter positive value if the field is in the positive z-direction and negative valu if the field is in the negative z-direction. E,Value Units Submit Request Answer Part C Find the electric field (magnitude and direction) at point on the r-axis at a -0.200 nm Express your answer with the appropriate units. Enter positive value if the field is in the positive a-direction and negative valu if the field is in the negative a-direction. E-Value Units
Answer #1

a) E = k Q / r2

E = 9 * 109 * [-3.5/0.22 + 6/(0.800-0.200)2] * 10-9

= -637.5 N/C

direction is towards left or negative x direction

b) E = k Q / r2

E = 9 * 109 * [-3.5/1.22 - 6/(1.2-0.8)2] * 10-9

= -359 N/C

direction is towards left or negative x direction

c) E = k Q / r2

E = 9 * 109 * [3.5/0.22 + 6/(0.8+0.2)2] * 10-9

= 841.5 N/C

direction is towards right or positive x direction